omework Consider the following time series data Month 1 2 3 4 5 6 7 Value 21 14 18 13 18 21 14 a. Which of the following is a correct time series plot for this data? や" -Select your answer- What type of pattern exists in the data? -select your answer- b. Develop the three-month moving average forecasts for this time series. Compute MSE and a forecast for month 8 (to 2 decimals if necessary). Enter negative values as negative number. Squared Forecast Forecast Time Series Month Value Forecast Error Error 2 14 3 18 4 13 5 18 6 21 7 14 Totals MSE The forecast for month 8 C. Use α-.2 to compute the exponential smoothing forecasts for the time series. Compute MSE and a forecast for month 8 (to 2 decimals). Enter negative values as negative number Time Squared Forecast Forecast Series Month Value Forecast Error Error 1 21 2 14 3 18
2 to compute the exponential smoothing forecasts for the me series Compute M SE and a forecast or 8 to 2 dec mals Ente negative values as negati e number c. Use α on Time Series Squared Forecast Forecast Month Value Forecast ErrorError 1 21 2 14 3 18 4 13 5 18 6 21 714 Totals MSE The forecast for month 8 d. Compare the three-month moving average approach with he exponential smoothing approach using α- 2. which appears to provide more accurate o ea ts base an M Select your answer e. Use a smoothing constant of α-,4 to compute the MSE. (to 2 decimals) Does a smoothing constant of .2 or .4 appear to provide more accurate forecasts based on MSE? -Select your answer-
d. Compare the three-month moving average approach with the exponential smoothing approach using α = .2, which appears to provide more accurate forecasts based on MSE? -Select your answer The three-month moving average provides a better forecast since it has a smaller MSE The exponential smoothing approach provides a better forecast since it has a smaller MSE The three-month moving average provides a better forecast since it has a larger MSE
e. Use a smoothing constant of α-.4 to compute the MSE. (to 2 decimals) Does a smoothing constant of .2 or .4 appear to provide more accurate forecasts based on MSE? -Select your answer - .2 .4
Answer #1

(a) right choice is first.

there is seasonality pattern in the data ( please give the options for appropriate answer)

25 20 15 10

(b)

t yt Forecast forecast error squared forecast error
1 21
2 14
3 18
4 13 17.6667 -4.6667 21.7778
5 18 15.0000 3.0000 9.0000
6 21 16.3333 4.6667 21.7778
7 14 17.3333 -3.3333 11.1111
total 63.6667

MSE=SQRT(63.6667/4)=3.9896

forecast for month 8=(18+21+14)/3=17.6667

(c)

alpha= 0.2
t Yt Ft=alpha*Y(t-1)+(1-alpha)*F(t-1) forecast error squared forecast error
1 21
2 14 21.0000 -7.0000 49.0000
3 18 19.6000 -1.6000 2.5600
4 13 19.2800 -6.2800 39.4384
5 18 18.0240 -0.0240 0.0006
6 21 18.0192 2.9808 8.8852
7 14 18.6154 -4.6154 21.3015
8 17.6923 total= 121.1857
9 MSE= 4.4942

MSE=sqrt(121.1857/6)=4.4942

forecast for the month 8=17.6923

(d) right choice is first.

since the 3-months moving average MSE is less than exponential smoothing (alpha=0.2 )

(e) alpha=0.4 is better ( than alpha=0.2).

alpha= 0.4
t Yt Ft=alpha*Y(t-1)+(1-alpha)*F(t-1) forecast error squared forecast error
1 21
2 14 21.0000 -7.0000 49.0000
3 18 18.2000 -0.2000 0.0400
4 13 18.1200 -5.1200 26.2144
5 18 16.0720 1.9280 3.7172
6 21 16.8432 4.1568 17.2790
7 14 18.5059 -4.5059 20.3033
8 16.7036 total= 116.5539
9 MSE= 4.4075